OpenStudy (anonymous):

an automobile traveling 95 km/h overtakes a 1.00km long train travelling in the same direction on a track parallel to the road. if the train's speed is 75km/h, how long does it take the car to pass it and how far will the car have traveled in this time? what are the results if the car and train are traveling in opposite directions?

OpenStudy (anonymous):

The velocity difference between the car and train is 20km/h t=(1km)/(20km/h)=0,05h t=3 minutes. when the opposite direction. The velocity difference is 170km/h t=(1km)/(170km/h)=5,88*10^-3 h t=21,17 seconds.

OpenStudy (anonymous):

How far will the car have traveled in 3.3 mins? (the length of the train is 1.10 km in this particular question) Are there unit conversions involved between hours and minutes?